OpenStudy (anonymous):

identify one of the factors of 32r^2+24ry-4xr-3xy

OpenStudy (disco619):

8r(4^2+3y) i can't do more.... CONFUSING!!!

OpenStudy (disco619):

I don't know if what i did till now is right or not...

OpenStudy (anonymous):

alright thanks though and are ther more people in the morning

OpenStudy (anonymous):

since its pretty late

OpenStudy (anonymous):

to help out

OpenStudy (callisto):

32r^2+24ry-4xr-3xy Let's factorize the first two terms: 32r^2+24ry ^ What are the common factors of these two terms?

OpenStudy (anonymous):

i wanna say 2r,r

OpenStudy (disco619):

i'd say '8' and 'r' Take the Highest Multiple... I'll leave the Rest to Callisto

OpenStudy (callisto):

Let's say, we pick 2r out: 32r^2+24ry = 2r(16r+12y) It seems we can further factorize (16r+12y). What is the common factor in these two terms?

OpenStudy (anonymous):

4,2,1

OpenStudy (anonymous):

right ?

OpenStudy (callisto):

Take the largest one :) If you take 2, then we can further factorize it again. At last, we'll get 4 :D So, we pick 4 out. 32r^2+24ry = 2r(16r+12y) = 2r [ 4 (4r + 3y) ] = 8r (4r+3y) So far so good?

OpenStudy (anonymous):

yea is that the answer though ?

OpenStudy (callisto):

No. Now, factorize the last two terms: What are the common factors of -4xr-3xy?

OpenStudy (anonymous):

i got -1, and -x

OpenStudy (callisto):

Which one should you pick?

OpenStudy (anonymous):

umm -1 ?

OpenStudy (callisto):

Why?

OpenStudy (anonymous):

wait but its negative.

OpenStudy (anonymous):

thex could be before riight.

OpenStudy (callisto):

You can take out both -1 and x

OpenStudy (callisto):

Now, it's your turn to factorize -4xr-3xy. Can you give it a try?

OpenStudy (anonymous):

i got -x(4r+3y)

OpenStudy (callisto):

Right. So, for 32r^2+24ry-4xr-3xy, we get: 32r^2+24ry-4xr-3xy = 8r (4r+3y) -x(4r+3y) Is it clear to you?

OpenStudy (anonymous):

yess so thats the final answer or no lol.

OpenStudy (callisto):

No, one more step.

OpenStudy (anonymous):

okie dokie

OpenStudy (anonymous):

factor the last one ?

OpenStudy (callisto):

Factorize this: 8r (4r+3y) -x(4r+3y)

OpenStudy (anonymous):

i got (4r+3y) and (8r-x)

OpenStudy (callisto):

When you factorize by grouping, you make it into factored form..So, it is (8r-x)(4r+3y)

OpenStudy (anonymous):

and my choices are 1. (x-8r) 2. (8x-r) 3. (4r+3y) 4. none and 5. (3r+4y)

OpenStudy (anonymous):

so i choose 3 right ?

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

Wow your an expert thank you so much.

OpenStudy (callisto):

You're welcome :)